OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y= 3/2x y=-1/2x +4 (___,___)

OpenStudy (anonymous):

I'm not going to be much help sorry...

OpenStudy (anonymous):

@helpout3 @ganeshie8 @WolfMaster @kirbykirby

OpenStudy (anonymous):

but

OpenStudy (anonymous):

u can google this

OpenStudy (anonymous):

okay, Thanks for trying :) @jakes4444444444444

OpenStudy (kirbykirby):

Since \(y\) is equal to \(\dfrac{3}{2}x\) and \(y\) is also equal to \(-\dfrac{1}{2}x+4\) Then you can just equate the 2 equations together, since they're both equal: \[ -\dfrac{1}{2}x+4=\frac{3}{2}x\] Add \(\dfrac{1}{2}x\) to both sides: \[ 4=2x\] Now you can solve for x. Once you find the value of x, plug in that value of x into any of your 2 initial equations (perhaps \(y=\dfrac{3}{2}x\) will be easier), and this will give you the value of y :)